OpenStudy (anonymous):

Find the solution of the equation: 3x squared + 20x +1= 8

OpenStudy (anonymous):

It doesn't look like you can factor, so you'll have to set the equation equal to zero (subtract 8 from both sides) and use the Quadratic Formula. (It's hard to type out, so here's a picture. http://www.crewtonramoneshouseofmath.com/images/quad_eq.gif) A is the x squared coefficient, b is the x coefficient, and c is the constant.

OpenStudy (anonymous):

Oops, that link doesn't work - here. http://www.scientificpages.net/math/images/Quadratic_Formula.gif

OpenStudy (anonymous):

that is not a formula that i am familiar with srry

OpenStudy (anonymous):

So, you look at your equation - 3x squared - 20x -7. The numbers you care about for the formula are 3, -20, and -7, because they are the numbers that correspond to a, b, and c in the formula (This only works if the equation is equal to zero). So you just plug in all the numbers in the right letter places in the formula. The plus and minus means there will be two answers - one when you subtract the stuff under the square root sign, and one where you add, before dividing.

OpenStudy (anonymous):

what i have is 3x squared + 20x -7=0 then 3x squared = 20x-1x-7 what do i do now

OpenStudy (anonymous):

Oh, +20x, sorry. Um, I think you pretty much have to use the quadratic formula. I'd be surprised if you hadn't talked about it in class, as I remember it was the first way I learned to solve quadratic equations (equations with x squared in them). So if it doesn't make sense, I'm not sure if there's anything I can do. When you plug in the numbers, you should get x= -(20) plus or minus square root of (20 squared - 4*3*(-7), all divided by 2*3, because, to reiterate, a=3, b=20, and c=-7. (So there are two answers, one where you add the square root stuff and one where you subtract it.) I can't help much more, sorry, because I came on here to review some trig for my calc test tomorrow and ended up seeing these questions and being like... "hey, I can answer that." But now I'm not studying...

OpenStudy (anonymous):

thank you for your time i have a weird teacher that teaches us different stuff than other people

OpenStudy (anonymous):

Sure, haha. It's possible I'm not seeing another way to do it. It's been a while. Good luck!

OpenStudy (anonymous):

thanks i will need it